OpenStudy (anonymous):

Evaluate. 3 (5x - 9)

OpenStudy (mimi_x3):

Evaluate? Do you mean expand ? 15x - 27

OpenStudy (anonymous):

\[3(5x-9)=15x-27\] as \[a(b-c)=ab-ac\]

OpenStudy (vishweshshrimali5):

15x - 27